OpenStudy (anonymous):

how do you parametrize 4x^2+y^2=4

OpenStudy (allank):

Oh nice, parametrization. First, do you know what we mean by parametrizing a figure?

OpenStudy (anonymous):

well its an ellipse, i need to parametrize it so i can use it with stokes theorem

OpenStudy (kenljw):

To normalize set all coefficients to one 4x^2+y^2=4 x^2 + (y/2)^2 = 1

OpenStudy (jhannybean):

solve for one, substitute it for the other?

OpenStudy (anonymous):

KenLJW ive already done that, i need to get it into the form x=rcost y=rsint, but that 4 messes everything up.

OpenStudy (allank):

Alright, the main idea behind parametrization is obtaining a function whose image is the figure...in this case an ellipse (you can read about that later) The parametrization I think you want is for R1 -> R2 Where you have a parametrization that takes in one variable and outputs values on the ellipse. For now, I can only think of how to get the parametrization of the top half of the ellipse. We have 4x^2+y^2=4 Solving for y in terms of x y=sqrt(4-4x^2) Let's create a parametrization x, with variable t. x(t)=(t,sqrt(4-4t^2)) Plugging in different t values yields the upper half of the ellipse.

OpenStudy (kenljw):

for unit circle x^2 + y'^2 =1 y' = y/2 sin = y' = y/2 cos = x

OpenStudy (kenljw):

not sure

OpenStudy (anonymous):

ok the answer given, is "4x^2 + y 2 = 4 or x^2 + (y^2)/4 = 1, so x = r cos q and y = 2r sin q" can someone show me how they got to this answer ?

OpenStudy (kenljw):

That's what I had r = 1, angle q

OpenStudy (anonymous):

hmm, im still quite confused. thanks for trying though

OpenStudy (kenljw):

x^2 + y'^2 = r y' =y/2 soh cah toa sin(q) = y'/r =y/2r therefore y =2r sin(q)